-TOPIC_ inversion of sugar Cane
I plotted a graph of ln(Optical rotation -
Infinite optical rotation) vs time..obtained the pseudo rate constant
How do I use optical rotation to determine the rotation at time zero
1
Expert's answer
2019-11-11T07:47:30-0500

Inversion of cane sugar is a pseudo first order reaction.

"Rate constant (k)×t= 2.303log((a_t-a_{\infty})/(a_0-a_{\infty}))"

"Rate constant (k)×t= 2.303log(a_t-a_{\infty})-2.303log(a_0-a_{\infty})"

"t= 2.303log(a_t-a_{\infty})/k-2.303log(a_0-a_{\infty})/k"

at= optical rotation at time "t=t"

a0= optical rotation at time "t=0"

"a_{\infty}=optical" rotation after infinite time

It is of the form "x=my+c"

"x=t"

"m=2.303\/k"

"y=log(a_t-a_{\infty})"

"c=-(2.303log(a_0-a_{\infty}))/k"

Now in the graph, when "x=0" ,we obtain x intercept which is equal to c.

We know k,"a_{\infty}" and c from the graph so using all these in equation (1) we obtain optical rotation at time 0.